# SDEs with supercritical distributional drifts

### v1

@@ 5.2. Generalized martingale problems
${\mathbb{E}}^{{\mathbb{P}}_{i}}\left(\int_{t_{n}}^{T}f(s,{w}_{s}){\mathord{{\rm d}}}s\,g({w}_{t_{1}}\cdots,{w}_{t_{n}})\right)={\mathbb{E}}^{{\mathbb{P}}_{i}}\Big{(}{\mathcal{S}}_{f}(t_{n},{w}_{t_{n}})g({w}_{t_{1}}\cdots,{w}_{t_{n}})\Big{)},\quad i=1,2,$
@@
Let $n_{k}$ be any subsequence and $n_{k}^{\prime}$ be the subsubsequence of $n_{k}$ used in the definition of ${\mathcal{S}}_{f}$. By Theorems 5.10 and 5.7, any accumulation point of ${\mathbb{P}}_{n^{\prime}_{k}}$ is a generalized martingale solution, and meanwhile, a weak solution. By the uniqueness of generalized martingale solutions, for any initial value $\mu$, without further selecting a subsequence, ${\mathbb{P}}_{n^{\prime}_{k}}$ weakly converges to the weak solution of SDE (1.1). Finally, the Markov property follows by the uniqueness of generalized martingale solutions (see [12]). If $b=b_{1}+b_{2}$, where $b_{1}\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}$ and $b_{2}\in{\mathbb{L}}^{2}_{T}L^{2}$, then PDE (5.16) has a unique weak solution. Thus, no necessary to select a subsequence, ${\mathcal{S}}_{f}$ is well-defined, and as above, the whole sequence ${\mathbb{P}}_{n}$ weakly converges to the weak solution of SDE (1.1). ∎
@@ 6.1. Vector-valued Gaussian fields
    For each Borel measurable set $A\subset{\mathbb{R}}^{d}$, $(\mu_{ij}(A))_{i,j=1}^{m}$ is a symmetric semi-positive definite Hermit matrix, that is, for any complex numbers $(a_{i})^{m}_{i=1}$,
@@
In general, the element in ${\mathbb{H}}_{0}$ may be not a Schwartz distribution.
@@
by the dominant convergence theorem, we have
@@
The continuity of $x\mapsto U_{\varphi}(x)$ now follows by Lemma 6.2. Similarly, by (6.3), we have (6.6). Moreover, by the the hypercontractivity of Gaussian random variables, we have for any $p\geqslant 2$,
@@
Below, for a Banach space ${\mathbb{B}}$, let ${\mathbf{B}}^{s}_{p}({\mathbb{B}}):={\mathbf{B}}^{s}_{p,\infty}({\mathbb{B}})$ be the ${\mathbb{B}}$-valued Besov space, and for any $\rho_{\kappa}(x)=(1+|x|)^{-\kappa}$, where $\kappa\in{\mathbb{R}}$, let ${\mathbf{B}}^{s}_{p}(\rho_{\kappa};{\mathbb{B}})$ be the weighted Besov space, which is defined by (see [22])
$\|f\|_{{\mathbf{B}}^{s}_{p}(\rho_{\kappa};{\mathbb{B}})}:=\|f\rho_{k}\|_{{\mathbf{B}}^{s}_{p}({\mathbb{B}})}.$
@@
$|\phi_{j}(\xi)|\lesssim_{C}1\wedge\big{(}2^{\gamma j}(1+|\xi|)^{-\gamma}\big{)},\ \ j\geqslant-1,\ \ \xi\in{\mathbb{R}}^{d}.$ (6.9)
@@
For each $i=1,\cdots,m$, under (6.2), by definition (6.5) and the hypercontractivity of Gaussian random variables, we have for any $p\in[1,\infty)$ and $j=0,1,\cdots,$,
@@ 6.2. Examples: Divergence-free Gaussian field
Obviously, (6.2) holds. In this case, for $\gamma>0$, it is well known that for some $c_{d,\gamma}>0$ (see [37, p117, Lemma 2]),
@@
It is easy to see that (6.2) holds for $\ell<\gamma-d$, and for $a:=(a_{i})_{i=1}^{d}\subset{\mathbb{C}}$ and $\xi\in{\mathbb{R}}^{d}$,
@@
Thus (A) holds. Let $b^{(\gamma)}$ be the ${\mathbb{R}}^{d}$-valued Gaussian random field with matrix-valued spectral measure $\mu^{(\gamma)}({\mathord{{\rm d}}}x)$. Then by (6.8), for any $p\in[1,\infty)$, $\kappa>d/p$ and $\beta<\frac{\gamma-d}{2}$,
$b^{(\gamma)}\in L^{p}(\Omega;{\mathbf{B}}^{\beta}_{p}(\rho_{\kappa};{\mathbb{R}}^{m})).$
In particular, there is a ${\mathbf{P}}$-null set $N\subset\Omega$ so that for all $\omega\not\in N$,
$b^{(\gamma)}(\omega,\cdot)\in{\mathbf{B}}^{\beta}_{p}(\rho_{\kappa};{\mathbb{R}}^{m}),$
@@
where *** stands for the transpose of a row vector. In particular, one can apply Theorem 1.2 to solve the following SDE with $\gamma\in(d-2,d)$:
@@
However, ${\mathscr{S}}_{0}({\mathbb{R}}^{2})$ is not a deterministic class since for $f\in{\mathscr{S}}_{0}({\mathbb{R}}^{2})$, $\int_{{\mathbb{R}}^{2}}f(x){\mathord{{\rm d}}}x=0$. So we can not claim $b^{(0)}=\nabla^{\perp}(-\Delta)^{-\frac{1}{2}}U$ as distributions.
Example 6.11.
Let $b_{\varepsilon}$ be the Gaussian field with matrix-valued spectral measure (see Remark 1.6)
$\mu_{\varepsilon}({\mathord{{\rm d}}}\xi):={\mathbf{1}}_{|\xi|<1/\varepsilon}\Big{(}{\mathbb{I}}_{d\times d}-\frac{\xi\otimes\xi}{|\xi|^{2}}\Big{)}{\mathord{{\rm d}}}\xi.$
By Remark 6.1 and Theorem 6.6, for $p\in[1,\infty)$ and $\kappa>d/p$, we have
$b_{\varepsilon}(x):=b_{\varepsilon}(\tau_{x}\delta_{0})\in\cap_{\beta>0}L^{p}(\Omega;{\mathbf{B}}^{\beta}_{p}(\rho_{\kappa};{\mathbb{R}}^{m})).$
Let $20<\varepsilon<\varepsilon^{\prime}\leqslant 1/$ 2. Noting that
$({\mathbb{I}}-\Delta)^{-\frac{\beta}{2}}(b_{\varepsilon}-b_{\varepsilon^{\prime}})(x)=(b_{\varepsilon}-b_{\varepsilon^{\prime}})(\tau_{x}({\mathbb{I}}-\Delta)^{-\frac{\beta}{2}}\delta_{0}),$
for $\beta>1$ and $p\geqslant 2$, we have for all $x\in{\mathbb{R}}^{d}$,
$\|({\mathbb{I}}-\Delta)^{-\frac{\beta}{2}}(b_{\varepsilon}-b_{\varepsilon^{\prime}})(x)\|_{L^{p}(\Omega)}$ $\lesssim\|({\mathbb{I}}-\Delta)^{-\frac{\beta}{2}}(b_{\varepsilon}-b_{\varepsilon^{\prime}})(x)\|_{L^{2}(\Omega)}$
$=\left(\int_{{\mathbb{R}}^{2}}{\mathbf{1}}_{1/\varepsilon^{\prime}|\leqslant\xi|<1/\varepsilon}(1+|\xi|^{2})^{-\frac{\beta}{2}}{\mathord{{\rm d}}}\xi\right)^{\frac{1}{2}}\lesssim|\varepsilon-\varepsilon^{\prime}|^{\frac{\beta-1}{2}}.$
Therefore, as in Theorem 6.6, for any $p\in[1,\infty)$, $\kappa>d/p$ and $\beta>1$, there exists a vector field $b\in L^{p}(\Omega,{\mathbf{H}}^{-\beta}_{p}(\rho_{\kappa}))$ with $\mathord{{\rm div}}b=0$ and such that
$\lim_{\varepsilon\to 0}{\mathbf{E}}\|\rho_{\kappa}(b_{\varepsilon}-b)\|_{{\mathbf{H}}^{-\beta}_{p}}^{p}=0.$
On the other hand, fix $p\in[2,\infty)$ and $\kappa>d/p$. It is easy to see that
$\sup_{x\in{\mathbb{R}}^{d}}\|({\mathbb{I}}-\Delta)^{-\frac{1}{2}}b_{\varepsilon}(x)\|_{L^{p}(\Omega)}\leqslant C\sqrt{\ln(1/\varepsilon)},$
and so,
$\sup_{\varepsilon\in(0,1/2)}{\mathbf{E}}\left(\frac{\|\rho_{\kappa}b_{\varepsilon}\|_{{\mathbf{H}}^{-1}_{p}}}{\sqrt{\ln(1/\varepsilon)}}\right)^{p}=0.$
By a tightness argument and Skorokhod’s representation theorem, there is a subsequence $\varepsilon_{n}\to 0$, a probability space $(\widetilde{\Omega},\widetilde{{\mathscr{F}}},\widetilde{{\mathbb{P}}})$ and a family of Gaussian process $\{\widetilde{b}_{n}\}_{n=1}^{\infty}$ thereon such that
~$\widetilde{\mathbb{P}}\circ\widetilde{b}_{n}^{-1}={\mathbf{P}}\circ\left(\frac{{b}_{\varepsilon_{n}}}{\sqrt{\ln(1/\varepsilon_{n})}}\right)^{-1}\ \mbox{ in ${\mathbf{H}}^{-1}_{p}(\rho_{\kappa})$,}$
and, there is a $\widetilde{{\mathbb{P}}}$-null set $N$ such that for all $\widetilde{\omega}\notin N$,
$\sup_{n\in{\mathbb{N}}}\|\widetilde{b}_{n}(\widetilde{\omega},\cdot)\|_{{\mathbf{H}}^{-1}_{p}(\rho_{\kappa})}<\infty.$ (6.14)
@@ 6.3. Stochastic heat equations
By Duhamael’s formula, the solution is determined by the following formula:
@@
In particular, we can use Theorem 1.2 to time-dependent vector field $b^{(\gamma)}(t,x)=u^{(\gamma)}(t,x)$ with $\gamma\in(d-4,d-2)$.
Acknowledgement: We would like to thank Rongchan Zhu and Xiangchan Zhu for their invaluable contributions, particularly for their insights into the background of the critical cases and the uniqueness of generalized martingale solutions. This work is partially supported by NNSFC grants of China (Nos. 12131019), and the German Research Foundation (DFG) through the Collaborative Research Centre(CRC) 1283 “Taming uncertainty and profiting from randomness and low regularity in analysis, stochastics and their applications”.

### v3

@@ 5.2. Generalized martingale problems
${\mathbb{E}}^{{\mathbb{P}}_{i}}\left(\int_{t_{n}}^{T}f(s,{w}_{s}){\mathord{{\rm d}}}s\,g({w}_{t_{1}}\cdots,{w}_{t_{n}})\right)=-{\mathbb{E}}^{{\mathbb{P}}_{i}}\Big{(}{\mathcal{S}}_{f}(t_{n},{w}_{t_{n}})g({w}_{t_{1}}\cdots,{w}_{t_{n}})\Big{)},\quad i=1,2,$ (5.28)
@@
Let $n_{k}$ be any subsequence and $n_{k}^{\prime}$ be the subsubsequence of $n_{k}$ used in the definition of ${\mathcal{S}}_{f}$. By Theorems 5.10 and 5.7, any accumulation point of ${\mathbb{P}}_{n^{\prime}_{k}}$ is a generalized martingale solution, and meanwhile, a weak solution. By the uniqueness of generalized martingale solutions, for any initial value $\mu$, without further selecting a subsequence, ${\mathbb{P}}_{n^{\prime}_{k}}$ weakly converges to the weak solution of SDE (1.1). If $b=b_{1}+b_{2}$, where $b_{1}\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}$ and $b_{2}\in{\mathbb{L}}^{2}_{T}L^{2}$, then PDE (5.18) has a unique weak solution. Thus, no necessary to select a subsequence, ${\mathcal{S}}_{f}$ is well-defined, and as above, the whole sequence ${\mathbb{P}}_{n}$ weakly converges to the weak solution of SDE (1.1).
[…]
${\mathbb{E}}^{{\mathbb{P}}_{\mu}}\Big{[}M^{f}_{T}-M^{f}_{s}\big{|}{\mathscr{B}}_{s}\Big{]}=0{={\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[{\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[M^{f}_{T}-M^{f}_{s}\big{|}{\mathscr{B}}_{s}\right]\big{|}w_{s}\right]}={\mathbb{E}}^{{\mathbb{P}}_{\mu}}\Big{[}M^{f}_{T}-M^{f}_{s}\big{|}w_{s}\Big{]},$
which in turn implies by ${\mathcal{S}}_{f}(T)=0$ that
${\mathbb{E}}^{{\mathbb{P}}_{\mu}}[{\mathcal{S}}_{f}(s,w_{s})|{\mathscr{B}}_{s}]+{\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[\int_{s}^{T}f(r,w_{r}){\mathord{{\rm d}}}r\Big{|}{\mathscr{B}}_{s}\right]={\mathbb{E}}^{{\mathbb{P}}_{\mu}}[{\mathcal{S}}_{f}(s,w_{s})|w_{s}]+{\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[\int_{s}^{T}f(r,w_{r}){\mathord{{\rm d}}}r\Big{|}w_{s}\right].$
[…]
${\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[\int_{s}^{T}f(r,w_{r}){\mathord{{\rm d}}}r\Big{|}{\mathscr{B}}_{s}\right]={\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[\int_{s}^{T}f(r,w_{r}){\mathord{{\rm d}}}r\Big{|}w_{s}\right].$
In particular, for fixed $t\in(s,T]$ and $f\in C_{c}({\mathbb{R}}^{d})$, by mollifying approximation, we further have
${\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[\int_{(t-1/n)\vee s}^{t}f(w_{r}){\mathord{{\rm d}}}r\Big{|}{\mathscr{B}}_{s}\right]={\mathbb{E}}^{{\mathbb{P}}_{\mu}}\left[\int_{(t-1/n)\vee s}^{t}f(w_{r}){\mathord{{\rm d}}}r\Big{|}w_{s}\right],\ \ n\in{\mathbb{N}}.$
Multiplying both sides by $n$ and letting $n\to\infty$, by the dominated convergence theorem, we obtain
${\mathbb{E}}^{{\mathbb{P}}_{\mu}}[f(w_{t})|{\mathscr{B}}_{s}]={\mathbb{E}}^{{\mathbb{P}}_{\mu}}[f(w_{t})|{w_{s}}].$
The proof is complete. ∎
@@ 6.1. Vector-valued Gaussian fields
    For each Borel measurable set $A\subset{\mathbb{R}}^{d}$, $(\mu_{ij}(A))_{i,j=1}^{m}$ is a symmetric semi-positive definite Hermitian matrix, that is, for any complex numbers $(a_{i})^{m}_{i=1}$,
@@
In general, elements in ${\mathbb{H}}_{0}$ might not be Schwartz distribution.
@@
by the dominated convergence theorem, we have
@@
The continuity of $x\mapsto U_{\varphi}(x)$ now follows by Lemma 6.2. Similarly, by (6.3), we have (6.6). Moreover, by the hypercontractivity of Gaussian random variables, we have for any $p\geqslant 2$,
@@
Below, for a Banach space ${\mathbb{B}}$, let ${\mathbf{B}}^{s}_{p}({\mathbb{B}}):={\mathbf{B}}^{s}_{p,\infty}({\mathbb{B}})$ be the ${\mathbb{B}}$-valued Besov space. For $\kappa\in{\mathbb{R}}$, let $\rho_{\kappa}(x):=(1+|x|)^{-\kappa}$. Let ${\mathbf{B}}^{s}_{p}(\rho_{\kappa};{\mathbb{B}})$ be the weighted Besov space, which is defined by (see [23])
$\|f\|_{{\mathbf{B}}^{s}_{p}(\rho_{\kappa};{\mathbb{B}})}:=\|f\rho_{k}\|_{{\mathbf{B}}^{s}_{p}({\mathbb{B}})}.$ (6.8)
@@
$|\phi_{j}(\xi)|\lesssim 1\wedge\big{(}2^{\gamma j}(1+|\xi|)^{-\gamma}\big{)},\ \ j\geqslant-1,\ \ \xi\in{\mathbb{R}}^{d}.$ (6.10)
@@
For each $i=1,\cdots,m$, under (6.2), by definition (6.5) and the hypercontractivity of Gaussian random variables, we have for any $p\in[1,\infty)$ and $j=0,1,\cdots$,
@@ 6.2. Examples: Divergence-free Gaussian field
Obviously, (6.2) holds. In this case, for $\gamma>0$, it is well known that for some $c_{d,\gamma}>0$ (see [40, p117, Lemma 2]),
@@
It is easy to see that (6.2) holds for $\ell<\gamma-d$, and for $a:=(a_{i})_{i=1}^{d}\subset{\mathbb{C}}$ and $\xi\in{\mathbb{R}}^{d}{\backslash\{0\}}$,
@@
Thus (A) holds. Let $b^{(\gamma)}$ be the ${\mathbb{R}}^{d}$-valued Gaussian random field with matrix-valued spectral measure $\mu^{(\gamma)}({\mathord{{\rm d}}}x)$. Then by (6.9), for any $p\in[1,\infty)$, $\kappa>\frac{d}{p}$ and $\beta<\frac{\gamma-d}{2}$,
$b^{(\gamma)}\in L^{p}(\Omega;{\mathbf{B}}^{\beta}_{p}(\rho_{\kappa};{\mathbb{R}}^{d})).$
In particular, for any $\kappa>0$ arbitrarily small and $\beta<\frac{\gamma-d}{2}$, one can choose $p$ large enough so that $\kappa>\frac{d}{p}$ and $\beta+\frac{d}{p}<\frac{\gamma-d}{2}$, and for ${\mathbf{P}}$-almost all $\omega$,
$b^{(\gamma)}(\omega,\cdot)\in{\mathbf{B}}^{\beta+d/p}_{p}(\rho_{\kappa};{\mathbb{R}}^{d})\subset{\mathbf{B}}^{\beta}_{\infty}(\rho_{\kappa};{\mathbb{R}}^{d})=:{\mathbf{C}}^{\beta}_{\kappa}({\mathbb{R}}^{d}),$ (6.14)
@@
where * stands for the transpose of a row vector.
In particular, when $\gamma\in(d-2,d)$, we can take $\beta>-1$. In this case, for ${\mathbf{P}}$-almost all $\omega$, by Lemma A.2 in appendix, there exists a decomposition $b^{(\gamma)}=b^{(\gamma)}_{1}+b^{(\gamma)}_{2}$ with
$b^{(\gamma)}_{1}{(\omega,\cdot)}\in{\mathbf{C}}^{\beta-\kappa}_{0}({\mathbb{R}}^{d})),\ \ \mathord{{\rm div}}b^{(\gamma)}_{1}{(\omega,\cdot)}\in{\mathbf{C}}^{\beta-\kappa}_{0}({\mathbb{R}})\quad\text{and}\quad b^{(\gamma)}_{2}{(\omega,\cdot)}\in{\mathbf{C}}^{\kappa}_{-\beta+2\kappa}({\mathbb{R}}^{d}),$
where $\kappa\in(0,(1+\beta)/2)$. In particular, $b^{(\gamma)}_{2}{(\omega,\cdot)}$ is continuous and
$|b^{(\gamma)}_{2}({\omega,}x)|\lesssim(1+|x|)^{-\beta+2\kappa}\lesssim 1+|x|.$
Therefore, one can apply Theorem 1.2 to solve the following SDE
@@
However, ${\mathscr{S}}_{0}({\mathbb{R}}^{2})$ is not a determining class since for $f\in{\mathscr{S}}_{0}({\mathbb{R}}^{2})$, $\int_{{\mathbb{R}}^{2}}f(x){\mathord{{\rm d}}}x=0$. So we can not claim $b^{(0)}=\nabla^{\perp}(-\Delta)^{-\frac{1}{2}}U$ as distributions.
@@ 6.3. Stochastic heat equations
By Duhamel’s formula, the solution is determined by the following formula:
@@
In particular, we can use Theorem 1.2 and Lemma A.2, as in Example 6.8, to obtain the weak well-posedness of SDE with time-dependent vector field $b(t,x)=u^{(\gamma)}(t,x)$ with $\gamma\in(d-4,d-2)$.
Appendix A Auxiliary lemmas
In the appendix, we recall a useful decomposition result for a distribution $f\in{\mathbf{C}}^{\alpha}_{\kappa}$ that is taken from [21, Lemma 2.4]. Let $\kappa,\alpha\in{\mathbb{R}}$ and $m\in{\mathbb{N}}$. Recalling $\rho_{\kappa}(x)=(1+|x|)^{-\kappa}$, we introduce the weighted Hölder space by (see (6.8))
${\mathbf{C}}^{\alpha}_{\kappa}({\mathbb{R}}^{m}):=\left\{\begin{aligned} &{\mathbf{B}}^{\alpha}_{\infty}(\rho_{\kappa};{\mathbb{R}}^{m}),\ \alpha\notin{\mathbb{N}}_{0},\\ &{\mathbf{H}}^{\alpha}_{\infty}(\rho_{\kappa};{\mathbb{R}}^{m}),\ \alpha\in{\mathbb{N}}_{0}.\end{aligned}\right.$
For any $f\in{\mathbf{C}}^{\alpha}_{\kappa}({\mathbb{R}}^{m})$, we define
${\mathscr{U}}_{\geqslant}(f):=\sum_{k=-1}^{\infty}\phi_{k}({\mathbb{I}}-S_{k})f,\quad{\mathscr{U}}_{\leqslant}(f):=\sum_{k=-1}^{\infty}\phi_{k}S_{k}f,$
where $\phi_{k}$ is defined in (2.1) and $S_{k}$ is defined in (2.12). We shall use the following easy fact:
$\phi_{j}=\phi_{j}\tilde{\phi}_{j}\mbox{with }\tilde{\phi}_{j}:=\phi_{j-1}+\phi_{j}+\phi_{j+1}.$ (A.1)
We first introduce the following result.
Lemma A.1.
For any $\kappa,\alpha\in{\mathbb{R}}$, it holds that
$\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}\asymp\sup_{j\geqslant-1}\|\phi_{j}f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}.$ (A.2)
Proof.
First of all, by the definition and scaling of function $\phi_{j}$, for any fixed $m\in{\mathbb{N}}$,
$\sup_{j\geqslant-1}\|\phi_{j}\|_{{\mathbf{C}}^{m}_{0}}<\infty,$
which implies that for $m\geqslant|\alpha|$,
$\sup_{j\geqslant-1}\|\phi_{j}f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}{\lesssim}\sup_{j\geqslant-1}\|\phi_{j}\|_{{\mathbf{C}}^{m}_{0}}\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}\lesssim\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}.$
Thus it remains to show $\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}\lesssim\sup_{j\geqslant-1}\|\phi_{j}f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}.$
By interpolation, without loss of generality, we may assume $\alpha\in{\mathbb{Z}}$. Let $\alpha\in{\mathbb{N}}_{0}$. Thanks to $\mathrm{supp}\,\phi_{j}\subset B_{2^{j+2}/3}\setminus B_{2^{j-1}}$,
$\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}=\Big{\|}\sum_{j\geqslant-1}\phi_{j}f\Big{\|}_{{\mathbf{C}}^{\alpha}_{\kappa}}\lesssim\sup_{j\geqslant-1}\|\phi_{j}f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}.$
Thus we get (A.2) for $\alpha\in{\mathbb{N}}_{0}$. On the other hand, by duality and (A.1), we have
$\|f\|_{{\mathbf{C}}^{-\alpha}_{\kappa}}$ $=\|\rho_{\kappa}f\|_{{\mathbf{C}}^{-\alpha}}=\sup_{\|g\|_{{\mathbf{H}}^{\alpha}_{1}}\leqslant 1}|\langle\rho_{\kappa}f,g\rangle|=\sup_{\|g\|_{{\mathbf{H}}^{\alpha}_{1}}\leqslant 1}\Big{|}\sum_{j}\langle\rho_{\kappa}\tilde{\phi}_{j}f,{\phi}_{j}g\rangle\Big{|}$
$\leqslant\sup_{\|g\|_{{\mathbf{H}}^{\alpha}_{1}}\leqslant 1}\sum_{j}\Big{(}\|\phi_{j}g\|_{{\mathbf{H}}^{\alpha}_{1}}\|\rho_{\kappa}\tilde{\phi}_{j}f\|_{{\mathbf{C}}^{-\alpha}}\Big{)}\lesssim\sup_{\|g\|_{{\mathbf{H}}^{\alpha}_{1}}\leqslant 1}\sum_{j}\|\phi_{j}g\|_{{\mathbf{H}}^{\alpha}_{1}}\sup_{j}\|\phi_{j}f\|_{{\mathbf{C}}^{-\alpha}_{\kappa}}.$
We note that
$\sum_{j}\|\phi_{j}g\|_{{\mathbf{H}}^{\alpha}_{1}}$ $\lesssim\sum_{j}\sum_{k\leqslant\alpha}\|\nabla^{\alpha-k}\phi_{j}\nabla^{k}g\|_{1}\lesssim\sum_{j}\sum_{k\leqslant\alpha}\|\nabla^{\alpha-k}\phi_{j}\tilde{\phi}_{j}\nabla^{k}g\|_{1}$
${\lesssim}\sum_{j}\sum_{k\leqslant\alpha}2^{(k-\alpha)j}\|\tilde{\phi}_{j}\nabla^{k}g\|_{1}\lesssim\sum_{k\leqslant\alpha}\sum_{j}\|{\phi}_{j}\nabla^{k}g\|_{1}\lesssim\sum_{k\leqslant\alpha}\|\nabla^{k}g\|_{1}\lesssim\|g\|_{{\mathbf{H}}^{\alpha}_{1}}.$
Then combining the above two estimates, we get $\|f\|_{{\mathbf{C}}^{-\alpha}_{\kappa}}\lesssim\sup_{j}\|\phi_{j}f\|_{{\mathbf{C}}^{-\alpha}_{\kappa}}$. ∎
Then we have the following result.
Lemma A.2.
Let $\alpha,\kappa\in{\mathbb{R}}$ and $f\in{\mathbf{C}}^{\alpha}_{\kappa}({\mathbb{R}}^{d})$. Then $f={\mathscr{U}}_{\geqslant}(f)+{\mathscr{U}}_{\leqslant}(f)$ and for any $\beta,\gamma>0$,
$\|{\mathscr{U}}_{\geqslant}(f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\lesssim\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}},\quad\|\mathord{{\rm div}}{\mathscr{U}}_{\geqslant}(f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\lesssim\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa+1}}+\|\mathord{{\rm div}}f\|_{{\mathbf{C}}^{\alpha}_{\kappa}},$ (A.3)
and
$\|{\mathscr{U}}_{\leqslant}(f)\|_{{\mathbf{C}}^{\alpha+\gamma}_{\kappa+\gamma}}\lesssim\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}.$ (A.4)
Proof.
By (A.2), we have
$\|{\mathscr{U}}_{\geqslant}(f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\lesssim\sup_{j}\|\phi_{j}{\mathscr{U}}_{\geqslant}(f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\leqslant\sup_{j}\left\|\sum_{k=-1}^{\infty}{\phi}_{j}\phi_{k}({\mathbb{I}}-S_{k})f\right\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\leqslant\sup_{k}\left\|\phi_{k}({\mathbb{I}}-S_{k})f\right\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}},$
where for $m\geqslant|\alpha|$
$\left\|\phi_{k}({\mathbb{I}}-S_{k})f\right\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\lesssim\|\phi_{k}\|_{{\mathbf{C}}^{m}_{-\beta}}\|({\mathbb{I}}-S_{k})f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa}}{\lesssim}2^{\beta k}\sum_{i=k}^{\infty}\|{\mathcal{R}}_{i}f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa}}.$
It follows from [23, (2.22) and (2.14)] that
$\|{\mathcal{R}}_{i}f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa}}$ $\lesssim\sup_{j}2^{(\alpha-\beta)j}\|{\mathcal{R}}_{j}{\mathcal{R}}_{i}f\|_{L^{\infty}(\rho_{\kappa})}\lesssim\sup_{j\sim i}2^{(\alpha-\beta)j}\|{\mathcal{R}}_{j}{\mathcal{R}}_{i}f\|_{L^{\infty}(\rho_{\kappa})}$
$\lesssim 2^{(\alpha-\beta)i}\|{\mathcal{R}}_{i}f\|_{L^{\infty}(\rho_{\kappa})}\lesssim 2^{-\beta i}\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}},$
which implies that
$\|{\mathscr{U}}_{\geqslant}(f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}$ $\lesssim\sup_{k}2^{\beta k}\sum_{i=k}^{\infty}\|{\mathcal{R}}_{i}f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa}}\lesssim\sup_{k}2^{\beta k}\sum_{i=k}^{\infty}2^{-\beta i}\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}\lesssim\|f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}.$
Similarly, we have
$\|\mathord{{\rm div}}{\mathscr{U}}_{\geqslant}(f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}$ $\leqslant\left\|\sum_{k=-1}^{\infty}\nabla\phi_{k}\cdot({\mathbb{I}}-S_{k})f\right\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}+\|{\mathscr{U}}_{\geqslant}(\mathord{{\rm div}}f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}},$
where
$\|{\mathscr{U}}_{\geqslant}(\mathord{{\rm div}}f)\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}\lesssim\|\mathord{{\rm div}}f\|_{{\mathbf{C}}^{\alpha}_{\kappa}}$
and
$\left\|\sum_{k=-1}^{\infty}\nabla\phi_{k}\cdot({\mathbb{I}}-S_{k})f\right\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}$ $\lesssim\sup_{j}\left\|\sum_{k=-1}^{\infty}{\phi}_{j}\nabla\phi_{k}\cdot({\mathbb{I}}-S_{k})f\right\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa-\beta}}$
$\leqslant\sup_{k}\|\nabla\phi_{k}\|_{{\mathbf{C}}^{m}_{-\beta-1}}\|({\mathbb{I}}-S_{k})f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa+1}}$
${\lesssim}\sup_{k}2^{\beta k}\|({\mathbb{I}}-S_{k})f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa+1}}\lesssim\|f\|_{{\mathbf{C}}^{\alpha-\beta}_{\kappa+1}}.$
The estimate for (A.4) is the same. ∎
Acknowledgement: We would like to thank Rongchan Zhu and Xiangchan Zhu for their invaluable contributions, particularly for their insights into the background of the critical cases and the uniqueness of generalized martingale solutions. We are also deeply thankful to the referees for their meticulous review of the manuscript and their constructive suggestions.